# Strong coupling between a single photon and a photon pair

### v2

@@ Abstract
The realization of strong nonlinear coupling between single photons has been a long-standing goal in quantum optics and quantum information science, promising wide impact applications, such as all-optical deterministic quantum logic and single-photon frequency conversion Lukin2014NPh ; Vuletic2012Nature . Here, we report an experimental observation of the strong coupling between a single photon and a photon pair in an ultrastrongly-coupled circuit-QED system. This strong nonlinear interaction is realized by introducing a detuned flux qubit working as an effective coupler between two modes of a superconducting coplanar waveguide resonator. The ultrastrong light–matter interaction breaks the excitation number conservation, and an external flux bias breaks the parity conservation. The combined effect of the two enables the strong one–two-photon coupling Nori2017PRA . Quantum Rabi-like avoided crossing is resolved when tuning the two-photon resonance frequency of the first mode across the single-photon resonance frequency of the second mode. Within this new photonic regime, we observe the second harmonic generation for a mean photon number below one. Our results represent a key step towards a new regime of quantum nonlinear optics, where individual photons can deterministically and coherently interact with each other in the absence of any stimulating fields Nori2017PRA ; Bouwmeester2006PRL .
@@
Strong interactions between single photons enable numerous applications in quantum information processing and simulations Lukin2014NPh . While photons in a vacuum do not interact with each other, when they are confined in nonlinear optical media, the light–matter interaction results in an effective photon–photon interaction Scully1999 ; Agarwal2012 . However, the nonlinearity of conventional materials is negligibly weak at the single-photon power level, making realizing strong nonlinear coupling between single photons extremely hard Lukin2014NPh . So far, important nonlinear optical phenomena like second harmonic generation (SHG) or down-conversion (DC) occur only statistically and in the presence of many photons, so a more or less small fraction of photons undergo fusion or fission process. Taking advantage of the resonant interaction between cavity photons and atoms in cavity QED Kimble2005Nature , using electromagnetically induced transparency (EIT) Lukin2009PRL , or collective effects in atomic ensembles Vuletic2012Nature , and applying high-power stimulating fields Aumentado2011NP are the common ways to circumvent this difficulty and to bring photon–photon interactions close to the quantum limit. However, an experimental observation of the strong coupling between different photon-number states Lukin2014NPh ; Bouwmeester2006PRL in the absence of control fields with many photons, a cornerstone of quantum many-body nonlinear optics, is still elusive.
@@
The development of superconducting quantum circuits (e.g., Martinis1985PRL ; Nakamura1999Nature ; Mooij1999Science ; Martinis2002PRL ; You2007PRB ; Koch2007PRA ) brings in new possibilities to tackle this problem. Josephson-junction (JJ)-based artificial atoms have a large dipole moment (not limited by the fine structure constant), which, combined with the small mode volume of superconducting coplanar resonators, allows the very large photon–artificial-atom interactions You2011Nature ; Nori2017PhyRep ; Mooij2004Nature ; Schoelkopf2004Nature ; Oliver2020ARCMP . The largest reported single-photon–single-qubit interactions have entered the ultrastrong and even deep strong coupling regimes Gross2010NP ; Mooij2010PRL ; Lupascu2017NP ; Semba2017NP ; Semba2018PRL ; Chen2017PRA ; Wang2020PRApplied ; Wang2023NC ; Tomonaga2023arXiv ; Nori2019NRP ; Solano2019RMP ; Hanggi2009PRA ; Nori2010PRA ; Solano2010PRL . This remarkable property makes Josephson artificial atoms an ideal effective medium to realize strong nonlinear coupling between single photons Nori2017PRA ; Semba2022PRRe ; Brien2021PRL ; Schoelkopf2013Nature ; Manucharyan2022Nature .
@@
Figure 1: |||| Setup. a, Schematic of the device. A flux qubit embedded in a ${\lambda}/{2}$ coplanar waveguide resonator, working as a nonlinear coupler between two modes of the resonator. The dashed cyan lines represent the vacuum current distribution of the $n=1$ (${\lambda}/{2}$) and $n=2$ ($\lambda$) modes of the resonator ${\kappa}_{\mathrm{out}}=3{\kappa}_{\mathrm{in}}$, where ${\kappa}_{\mathrm{in(out)}}$ is the loss rate from the input (output) port of the resonator, b, Optical image of the flux qubit. c, Effective coupling mechanism between the bare states $|2,0,g\rangle$ and $|0,1,g\rangle$ via virtual transitions involving intermediate states of the circuit-QED system. Here the first two entries in the kets denote the number of photons in the first two modes of the resonator, and the third indicates the qubit state ($|g\rangle$ is the ground state). d, Scheme of the energy levels at the flux offset corresponding to the minimum gap of the $|2,0,g\rangle-|0,1,g\rangle$ anticrossing, resulting in an effective coupling between these two states. A key feature of this configuration is that, at the minimum anticrossing gap, all the transitions shown by the arrows have comparable large efficiency, with transition matrix elements $|{X}_{1,0}|\simeq|{X}_{+,1}|\simeq|{X}_{+,0}|\sim 1$ (see Supplementary Fig. 2). Moreover, the transition energy $\tilde{\omega}_{+}$ is almost equal to $2\tilde{\omega}_{1}$, so that the second harmonic generation (with only two initial photons) and degenerate down conversion (with only one intial photon) can occur with unprecedented efficiency.
@@
Here, we propose and demonstrate the resonant strong coupling of a two-photon state with a one-photon state in a circuit-QED multimode photonic device. In this system, the ultrastrong light–matter interaction breaks the excitation number conservation and enables higher-order processes via virtual photons, and an external flux bias breaks the parity conservation Nori2017PRA ; Wang2023NC . The combined effect of the two enables the strong one–two-photon coupling.
@@
We realized a superconducting circuit constituted by a coplanar waveguide resonator embedded with an ultrastrongly-coupled flux qubit Mooij1999Science ; You2007PRB ; Chen2017PRA ; Wang2020PRApplied ; Wang2023NC ; Yan2016NC to implement a multimode system with its two-photon state of the lowest-energy mode strongly coupled to the one-photon state of the second mode (see Methods for details). Throughout this work, we refer to the $n$th mode as the $n\lambda/2$-mode. The flux qubit is significantly detuned with respect to the resonant frequencies of the two lowest energy modes of the resonator (dispersive regime), and acts as an effective coupler Nori2017PRA ; Semba2022PRRe ; Brien2021PRL between the $n=1$ and $n=2$ modes in the absence of any stimulating fields (Fig. 1c, d). The nonlinear coupling enables a photon pair in the $n=1$ mode to coherently interact with a single photon in the $n=2$ mode, a hallmark of which is a spectrally resolved mode splitting. We experimentally observed this photon–photon quantum Rabi-like splitting, resulting from the spontaneous hybridization of one- and two-photon states. It indicates that our system operates in the single-photon strong nonlinear coupling regime Lukin2014NPh ; Bouwmeester2006PRL , without involving real qubit excitations or any external driving mechanism. As we show here, within this regime, the SHG and DC processes can occur even when the input signals provide photons to the resonator with mean photon numbers below one.
@@
Figure 2: |||| Quantum Rabi-like splitting. Quantum Rabi-like splitting between a single photon and a photon pair. The left part in each panel (negative flux offset) reports the measured spectra. The corresponding calculated spectra are reported for positive flux offset (right), using the parity symmetry with respect to the flux offset of the theoretical calculations. a, Measured and calculated transmission spectra of the one-photon line of the $n=1$ mode of the resonator as a function of the flux offset. b, Measured and calculated transmission spectra showing the excitation of both the two-photon state of the $n=1$ mode and the one-photon state of the $n=2$ mode of the resonator. An avoided-level crossing between these two lines is clearly visible. From the fitting, we obtain the Rabi frequency for the one–two-photon coupling $g_{\mathrm{eff}}/2\pi=59\ \mathrm{MHz}$, the loss rate due to the input–output ports $({\kappa}_{\mathrm{in}}+\kappa_{\mathrm{out}})/2\pi=2.6\ \mathrm{MHz}$, the internal loss rate $\kappa_{\mathrm{int}}/2\pi=10.4\ \mathrm{MHz}$ (total loss rate of the resonator $\kappa_{\mathrm{tot}}/2\pi=(\kappa_{\mathrm{in}}+\kappa_{\mathrm{out}}+\kappa_{\mathrm{int}})/2\pi=13\ \mathrm{MHz}$), the intrinsic loss rate of the qubit ${\kappa}_{q}/2\pi=200\ \mathrm{MHz}$, and the pure dephasing rate of the qubit ${\kappa}_{q,\mathrm{dep}}/2\pi=200\ \mathrm{MHz}$ (the last two have a very weak influence on the calculated spectra).
@@
The Hamiltonian of the flux qubit can be written in the basis of two states with persistent currents ±$\pm I_{p}$ flowing in opposite directions around the qubit loop as (setting $\hslash=1$) $H_{q}=-(\Delta\sigma_{x}+\varepsilon\sigma_{z})/2$, where $\sigma_{x,z}$ are Pauli matrices, $\Delta$ and $\varepsilon=2I_{p}\delta\Phi_{\rm ext}$ are the tunnel splitting and the energy bias between the two basis states determined by an adjustable external flux bias $\delta\Phi_{\rm ext}$. The parameters $\Delta/2\pi=12.3\ \mathrm{GHz}$ and $I_{p}=60\ \mathrm{nA}$ are estimated by numerically diagonalizing the three-junction flux qubit Hamiltonian You2007PRB ; Yan2016NC . The resulting transition frequency between the two energy eigenstates $|g\rangle$ and $|e\rangle$ of the flux qubit is ${\omega}_{q}=\sqrt{\Delta^{2}+\varepsilon^{2}}$. A schematic of the device and the optical image of the flux qubit are shown in Fig. 1. For the flux qubit ultrastrongly coupled to the resonantor, the system can be described by a generalized two-mode quantum Rabi Hamiltonian Nori2017PhyRep
@@
where ${\omega}_{n}$ are the resonance frequencies of the $n=1,2$ modes of the resonator, $a_{n}^{\dagger}$ and $a_{n}$ are the corresponding creation and annihilation operators, and $g_{\mathrm{1,2}}$ are the coupling strengths of the flux qubit with the $n=1,2$ modes of the resonator. Owing to the inhomogeneous transmission line geometry (see Fig. 1b), the higher-mode frequency $\omega_{2}$ is not exactly twice the  fundamental resonance frequency $\omega_{1}$. Moreover, because of the contribution of the inductance across the qubit loop ($\delta{\mathrm{\Phi}}_{\mathrm{ext}}$-dependent) to the total inductance of the resonator, the resonance frequencies of the resonator (${\omega}_{1,2}$) become V-shaped around the optimal point ${\omega}_{1,2}(\delta{\mathrm{\Phi}}_{\mathrm{ext}})={\omega}_{1,2}(0)/\sqrt{1-\beta_{1,2}\cos(\theta)|\delta{\mathrm{\Phi}}_{\mathrm{ext}}|}$, with $\tan(\theta)=\Delta/\varepsilon$ and $\beta_{1,2}$ being constants Semba2017NP . Hereafter, $\tilde{\omega}_{j}$ denotes the energy eigenvalues of the Hamiltonian in Eq. (1) with respect to the ground state.
The coupling with the flux qubit determines an effective strong coupling between a single-photon state in mode $2$ ($|0,1,g\rangle$) and the two-photon state of mode $1$ ($|2,0,g\rangle$). As shown in Fig. 1c, such a coupling results from higher-order processes involving virtual intermediate transitions enabled by the ultrastrong qubit–resonator interaction Nori2019NRP and, unlike many processes in conventional nonlinear optics, it does not require any external drives. Such an effective interaction can be written as
$V_{\mathrm{eff}}=g_{\mathrm{eff}}(|\widetilde{0,1}\rangle\langle\widetilde{2,0}|+|\widetilde{2,0}\rangle\langle\widetilde{0,1}|)\,,$ (2)
@@
when ${\omega}_{2}\simeq{2\omega}_{1}$. The effective coupling goes to zero for $\theta=\pi/2$, when the system displays parity symmetry. The effective Hamiltonian in Eq. (1) describes the simultaneous annihilation of two photons in the $n=1$ mode and the creation of one photon in the $n=2$ mode as well as the inverse process. The coupling with the flux qubit introduces further Lamb shifts on the resonance frequencies of the resonator.
Figure 3: |||| Second harmonic generation. a, The amplitude of the SHG versus the external flux bias $\delta{\mathrm{\Phi}}_{\mathrm{ext}}$ and two times of the signal frequency applied at the $n=1$ mode of the resonator. b, Theoretical calculation corresponding to the plot in a. The power applied corresponds to an average photon number ${\overline{n}}_{1}\simeq 0.25$ in the resonator. c, The amplitude of the SHG versus the average photon number ${\overline{n}}_{1}$ in the resonator. The signal frequency applied at the $n=1$ mode of the resonator is ${\omega}_{1}/2\pi=4.9\ \mathrm{GHz}$ (at ${\delta\mathrm{\Phi}}_{\mathrm{ext}}=-45~{}{\rm m\Phi_{0}}$). The red solid curve is the theoretical fit.
@@
The device is placed inside a dilution refrigerator to be cooled down to a temperature of 20 mK. Considering the relevant frequencies (${\omega}_{1}$around $5.0\ \mathrm{GHz}$, ${\omega}_{2}$around $10.0\ \mathrm{GHz}$), the system nearly stays in its ground state at such a low temperature. To see the coupling between the two modes of the resonator, we measured the transmission spectra by applying a weak probe tone with its frequency scanning across the lowest transition energies $\tilde{\omega}_{j}$ with $j=1,3$. The spectra are fitted with the numerically calculated excitation energies of the system Hamiltonian $H_{\mathrm{s}}$. The parameters are determined to be ${\omega}_{1}/2\pi=5.0/\sqrt{1-\beta_{1}\cos(\theta)|\delta{\mathrm{\Phi}}_{\mathrm{ext}}|}\ \mathrm{GHz}$, ${\omega}_{2}/2\pi=9.7/\sqrt{1-\beta_{2}\cos(\theta)|\delta{\mathrm{\Phi}}_{\mathrm{ext}}|}\ \mathrm{GHz}$, $g_{1}/2\pi=2.815\ \mathrm{GHz}$, and $g_{2}/2\pi=2.180\ \mathrm{GHz}$, with $\beta_{1}=0.775\ \mathrm{(\Phi_{0})^{-1}}$ and $\beta_{2}=0.919\ \mathrm{(\Phi_{0})^{-1}}$. The coupling strengths ${g_{1(2)}}/{{\omega}_{1(2)}}=0.563(0.225)>0.1$ at the optimal point (${\delta\mathrm{\Phi}}_{\mathrm{ext}}=0$) provide evidence our system is in the multimode ultrastrong coupling regime. Parameters from the fit are used as input for density matrix calculations which are then compared with the measured spectra. Theoretical spectra throughout this work are obtained by using a generalized master equation Petruccione2002Theory ; Settineri2018Dissipation , where the interaction with the environment takes into account the hybridization of the subsystems due to the ultrastrong coupling (see Methods for details).
@@
Figure 2 reports the measured (left) and calculated (right) transmission spectra versus the flux offset. The calculated spectra are proportional to the time derivative of the expectation value of the two-mode field momentum operator: $\hat{X}=i(\hat{a}_{1}-\hat{a}_{1}^{\dagger})+i\sqrt{\omega_{2}/\omega_{1}}(\hat{a}_{2}-\hat{a}_{2}^{\dagger})$ (see Supplemeantary Information for details). Figure 2a shows the lowest energy transition $\tilde{\omega}_{1}$ corresponding to the excitation of the state $|E_{1}\rangle\equiv|\widetilde{1,0}\rangle\simeq|1,0,g\rangle$. Figure 2b shows the transitions at $\tilde{\omega}_{2}$ and $\tilde{\omega}_{3}$. When the two-photon resonance frequency of the $n=1$ mode of the resonator gets across the single-photon resonance frequency of the $n=2$ mode ($\omega_{2}\simeq 2\omega_{1}$), a clear avoided crossing, induced by the effective interaction in Eq. (2), is observed in Fig. 2b, indicating that our system enters the single-photon strong nonlinear coupling regime Bouwmeester2006PRL $g_{\mathrm{eff}}>\kappa_{\mathrm{tot}}$, where ${\kappa}_{\mathrm{tot}}$ is the total loss rate of the resonator. The nonlinear coupling strength $g_{\mathrm{eff}}/2\pi$ extracted from the fit is 59 MHz (half the separation between the double peaks at the minimum splitting), which qualitatively agrees with the perturbative result given by Eq. 3 (47 MHz). While for flux offsets outside the two-modes resonance condition, the system eigenstates are approximately $|2,0,g\rangle$ and $|0,1,g\rangle$; at the anticrossing they are in good approximation their symmetric and antisymmetric superpositions $|\psi_{\pm}\rangle=|2,0,g\rangle\pm|0,1,g\rangle$. Note that outside the avoided-level crossing in Fig. 2b, the spectral line corresponding to the excitation of the two-photon state $|2,0,g\rangle$ is poorly excited by the weak probe tone. Two-photon contributions get significantly excited only owing to their hybridization with one-photon states at the avoided crossing. The emission of photon pairs under very weak excitation is maximized in this region (see Supplementary Fig. 3). Theoretical calculations accurately reproduce all the spectra in Fig. 2b, thus confirming our interpretation of the data. A peculiar feature of this extreme quantum nonlinear regime is that linear spectra in Fig. 2b are able to evidence quantum nonlinear optical processes.

### v3

# Strong coupling between a single-photon and a two-photon Fock state

@@ Abstract
The realization of strong nonlinear coupling between single photons has been a long-standing goal in quantum optics and quantum information science, promising wide impact applications, such as all-optical deterministic quantum logic and single-photon frequency conversion. Here, we report an experimental observation of the strong coupling between a single-photon and a two-photon Fock state in an ultrastrongly-coupled circuit-QED system. This strong nonlinear interaction is realized by introducing a detuned flux qubit working as an effective coupler between two modes of a superconducting coplanar waveguide resonator. The ultrastrong light–matter interaction breaks the excitation number conservation, and an external flux bias breaks the parity conservation. The combined effect of the two enables the strong one–two-photon coupling. Quantum Rabi-like avoided crossing is resolved when tuning the two-photon resonance frequency of the first mode across the single-photon resonance frequency of the second mode. Within this new photonic regime, we observe the thresholdless second harmonic generation for a mean photon number below one. Our results represent a key step towards a new regime of quantum nonlinear optics, where individual photons can deterministically and coherently interact with each other in the absence of any stimulating fields.
Introduction
@@
Strong interactions between single photons enable numerous applications in quantum information processing and simulations Lukin2014NPh . While photons in a vacuum do not interact with each other, when they are confined in nonlinear optical media, the light–matter interaction results in an effective photon–photon interaction Scully1999 ; Agarwal2012 . However, the nonlinearity of conventional materials is negligibly weak at the single-photon power level, making realizing strong nonlinear coupling between single photons extremely hard Lukin2014NPh .
So far, important nonlinear optical phenomena like second harmonic generation (SHG) or down-conversion (DC) occur only statistically and in the presence of many photons, so a more or less small fraction of photons undergo fusion or fission process. Taking advantage of the resonant interaction between cavity photons and atoms in cavity QED Kimble2005Nature , using electromagnetically induced transparency (EIT) Lukin2009PRL , or collective effects in atomic ensembles Vuletic2012Nature , and applying high-power stimulating fields Aumentado2011NP are the common ways to circumvent this difficulty and to bring photon–photon interactions close to the quantum limit. However, an experimental observation of the strong coupling between different photon-number states Lukin2014NPh ; Bouwmeester2006PRL in the absence of control fields with many photons, a cornerstone of quantum many-body nonlinear optics, is still elusive.
@@
The development of superconducting quantum circuits (e.g., Martinis1985PRL ; Nakamura1999Nature ; Mooij1999Science ; Martinis2002PRL ; You2007PRB ; Koch2007PRA ) brings in new possibilities to tackle this problem. Josephson-junction (JJ)-based artificial atoms have a large dipole moment (not limited by the fine structure constant), which, combined with the small mode volume of superconducting coplanar resonators, allows the very large photon–artificial-atom interactions You2011Nature ; Nori2017PhyRep ; Mooij2004Nature ; Schoelkopf2004Nature ; Oliver2020ARCMP . The largest reported single-photon–single-qubit interactions have entered the ultrastrong and even deep strong coupling regimes Gross2010NP ; Mooij2010PRL ; Lupascu2017NP ; Semba2017NP ; Semba2018PRL ; Chen2017PRA ; Wang2020PRApplied ; Wang2023NC ; Tomonaga2023arXiv ; Nori2019NRP ; Solano2019RMP ; Hanggi2009PRA ; Nori2010PRA ; Solano2010PRL . This remarkable property makes Josephson artificial atoms an ideal effective medium to realize strong nonlinear coupling between single photons Nori2017PRA ; Semba2022PRRe ; Brien2021PRL ; Schoelkopf2013Nature ; Manucharyan2022Nature ; Fraudet2024arXiv .
@@
Figure 1: $|$ Setup. a, Schematic of the device. A flux qubit embedded in a ${\lambda}/{2}$ coplanar waveguide resonator, working as a nonlinear coupler between two modes of the resonator. The dashed cyan lines represent the vacuum current distribution of the $n=1$ (${\lambda}/{2}$) and $n=2$ ($\lambda$) modes of the resonator ${\kappa}_{\mathrm{out}}=3{\kappa}_{\mathrm{in}}$, where ${\kappa}_{\mathrm{in(out)}}$ is the loss rate from the input (output) port of the resonator, b, Optical image of the flux qubit. c, Effective coupling mechanism between the bare states $|2,0,g\rangle$ and $|0,1,g\rangle$ via virtual transitions involving intermediate states of the circuit-QED system. Here the first two entries in the kets denote the number of photons in the first two modes of the resonator, and the third indicates the qubit state ($|g\rangle$ is the ground state). d, Scheme of the energy levels at the flux offset corresponding to the minimum gap of the $|2,0,g\rangle-|0,1,g\rangle$ anticrossing, resulting in an effective coupling between these two states. A key feature of this configuration is that, at the minimum anticrossing gap, all the transitions shown by the arrows have comparable large efficiency, with transition matrix elements $|{X}_{1,0}|\simeq|{X}_{+,1}|\simeq|{X}_{+,0}|\sim 1$ (see Supplementary Fig. 2). Moreover, the transition energy $\tilde{\omega}_{3}$ is almost equal to $2\tilde{\omega}_{1}$, so that the second harmonic generation (with only two initial photons) and degenerate down conversion (with only one initial photon) can occur efficiently at sub-photon input levels.
@@
Here, we propose and demonstrate the resonant strong coupling of a two-photon state with a one-photon state in a circuit-QED multimode photonic device. In this system, the ultrastrong light–matter interaction breaks the excitation number conservation and enables higher-order processes via virtual photons, and an external flux bias breaks the parity conservation Nori2017PRA ; Wang2023NC ; Deppe2008NP . The combined effect of the two enables the strong one–two-photon coupling.
@@ Results
We realized a superconducting circuit constituted by a coplanar waveguide resonator embedded with an ultrastrongly-coupled flux qubit Mooij1999Science ; You2007PRB ; Chen2017PRA ; Wang2020PRApplied ; Wang2023NC ; Yan2016NC to implement a multimode system with its two-photon state of the lowest-energy mode strongly coupled to the one-photon state of the second mode (see Methods for details). Throughout this work, we refer to the $n$th mode as the $n\lambda/2$-mode. The flux qubit is significantly detuned with respect to the resonant frequencies of the two lowest energy modes of the resonator (dispersive regime), and acts as an effective coupler Nori2017PRA ; Semba2022PRRe ; Brien2021PRL between the $n=1$ and $n=2$ modes in the absence of any stimulating fields (Fig. 1c, d). The nonlinear coupling enables two photons in the $n=1$ mode to coherently interact with a single photon in the $n=2$ mode, a hallmark of which is a spectrally resolved mode splitting. We experimentally observed this photon–photon quantum Rabi-like splitting, resulting from the spontaneous hybridization of one- and two-photon states. It indicates that our system operates in the single-photon strong nonlinear coupling regime Lukin2014NPh ; Bouwmeester2006PRL , without involving real qubit excitations or any external driving mechanism. As we show here, within this regime, the SHG and stimulated DC processes can occur even when the input signals provide photons to the resonator with mean photon numbers below one.
@@
Figure 2: $|$ Quantum Rabi-like splitting. Quantum Rabi-like splitting between a single-photon and a two-photon Fock state. The left part in each panel (negative flux offset) reports the measured spectra. The corresponding calculated spectra are reported for positive flux offset (right), using the parity symmetry with respect to the flux offset of the theoretical calculations. a, Measured and calculated transmission spectra of the one-photon line of the $n=1$ mode of the resonator as a function of the flux offset. b, Measured and calculated transmission spectra showing the excitation of both the two-photon state of the $n=1$ mode and the one-photon state of the $n=2$ mode of the resonator. An avoided-level crossing between these two lines is clearly visible. From the fitting, we obtain the Rabi frequency for the one–two-photon coupling $g_{\mathrm{eff}}/2\pi=59\ \mathrm{MHz}$, the loss rate due to the input–output ports $({\kappa}_{\mathrm{in}}+\kappa_{\mathrm{out}})/2\pi=2.6\ \mathrm{MHz}$, the internal loss rate $\kappa_{\mathrm{int}}/2\pi=10.4\ \mathrm{MHz}$ (total loss rate of the resonator $\kappa_{\mathrm{tot}}/2\pi=(\kappa_{\mathrm{in}}+\kappa_{\mathrm{out}}+\kappa_{\mathrm{int}})/2\pi=13\ \mathrm{MHz}$), the intrinsic loss rate of the qubit ${\kappa}_{q}/2\pi=200\ \mathrm{MHz}$, and the pure dephasing rate of the qubit ${\kappa}_{q,\mathrm{dep}}/2\pi=200\ \mathrm{MHz}$ (the last two have a very weak influence on the calculated spectra).
@@
The Hamiltonian of the flux qubit can be written in the basis of two states with persistent currents $\pm I_{p}$ flowing in opposite directions around the qubit loop as (setting $\hslash=1$) $H_{q}=-(\Delta\sigma_{x}+\varepsilon\sigma_{z})/2$, where $\sigma_{x,z}$ are Pauli matrices, $\Delta$ and $\varepsilon=2I_{p}\delta\Phi_{\rm ext}$ are the tunnel splitting and the energy bias between the two basis states determined by an adjustable external flux bias $\delta\Phi_{\rm ext}$. The parameters $\Delta/2\pi=12.3\ \mathrm{GHz}$ and $I_{p}=60\ \mathrm{nA}$ are estimated by numerically diagonalizing the three-junction flux qubit Hamiltonian You2007PRB ; Yan2016NC . The resulting transition frequency between the two energy eigenstates $|g\rangle$ and $|e\rangle$ of the flux qubit is ${\omega}_{q}=\sqrt{\Delta^{2}+\varepsilon^{2}}$. A schematic of the device and the optical image of the flux qubit are shown in Fig. 1.
For the flux qubit ultrastrongly coupled to the resonantor, the system can be described by a generalized two-mode quantum Rabi Hamiltonian Nori2017PhyRep
@@
where ${\omega}_{n}$ are the resonance frequencies of the $n=1,2$ modes of the resonator, $a_{n}^{\dagger}$ and $a_{n}$ are the corresponding creation and annihilation operators, and $g_{\mathrm{1,2}}$ are the coupling strengths of the flux qubit with the $n=1,2$ modes of the resonator.
In Supplementary Fig. 2, we analyze the influence of higher-energy modes, showing that the two-mode approximation is valid. Owing to the inhomogeneous transmission line geometry due to the qubit presence (see Fig. 1b), the higher-mode frequency $\omega_{2}$ is not exactly twice the fundamental resonance frequency $\omega_{1}$. Moreover, because of the contribution of the inductance across the qubit loop ($\delta{\mathrm{\Phi}}_{\mathrm{ext}}$-dependent) to the total inductance of the resonator, the resonance frequencies ${\omega}_{1,2}$ of the resonator become V-shaped around the optimal point
${\omega}_{1,2}(\delta{\mathrm{\Phi}}_{\mathrm{ext}})={\omega}_{1,2}(0)/\sqrt{1-\beta_{1,2}\cos(\theta)|\delta{\mathrm{\Phi}}_{\mathrm{ext}}|},$ (2)
with $\tan(\theta)=\Delta/\varepsilon$ and $\beta_{1,2}$ being constants Semba2017NP . Hereafter, $\tilde{\omega}_{j}$ denotes the energy eigenvalues of the Hamiltonian in Eq. (1) with respect to the ground state.
The coupling with the flux qubit determines an effective strong coupling between a single-photon state in mode 22 ($|0,1,g\rangle$) and the two-photon state of mode 11 ($|2,0,g\rangle$). As shown in Fig. 1c, such a coupling results from higher-order processes involving virtual intermediate transitions enabled by the ultrastrong qubit–resonator interaction Nori2019NRP and, unlike many processes in conventional nonlinear optics, it does not require any external drives. Such an effective interaction can be written as
[…]
@@
when ${\omega}_{2}\simeq{2\omega}_{1}$. The effective coupling goes to zero for $\theta=\pi/2$, when the system displays parity symmetry. The effective Hamiltonian in Eq. (3) describes the simultaneous annihilation of two photons in the $n=1$ mode and the creation of one photon in the $n=2$ mode as well as the inverse process. The coupling with the flux qubit introduces further Lamb shifts on the resonance frequencies of the resonator. The qubit population is very small at the one–two-photon avoided level crossing and can be traced out from the dynamics (see Supplementary Fig. 4).
Figure 3: $|$ Second harmonic generation. a, The amplitude of the SHG versus the external flux bias $\delta{\mathrm{\Phi}}_{\mathrm{ext}}$ and the SHG frequency. The efficiency $\eta\equiv S_{21}^{(2\omega)}/S_{21}^{(\omega)}$ of the SHG is about 0.1 at the point where the SHG amplitude is maximized (${\omega}_{1}/2\pi=4.9\ \mathrm{GHz}$ and ${\delta\mathrm{\Phi}}_{\mathrm{ext}}=-45~{\rm m\Phi_{0}}$), which qualitatively agrees with the theoretical result in Supplementary Fig. 2. b, Theoretical calculation corresponding to the plot in a. c, The amplitude of the SHG versus the average photon number ${\overline{n}}_{1}$ in the resonator. The signal frequency applied at the $n=1$ mode of the resonator is ${\omega}_{1}/2\pi=4.9\ \mathrm{GHz}$ (at ${\delta\mathrm{\Phi}}_{\mathrm{ext}}=-45~{\rm m\Phi_{0}}$). The red solid curve is the theoretical fit. The average photon number in c is determined by contrasting the experimental data—plotting the maximum second-harmonic generation (SHG) amplitude against input power—with the theoretical simulation outlined in Supplementary Eq. (35), as depicted in c. The power applied in a and b corresponds to an average photon number ${\overline{n}}_{1}\simeq 0.25$ in the resonator.
@@
The device is placed inside a dilution refrigerator to be cooled down to a temperature of 20 mK. Considering the relevant frequencies (${\omega}_{1}/2\pi\$around 5.0GHz5.0\ $\mathrm${GHz}, ${\omega}_{2}/2\pi\$around 10.0GHz10.0\ $\mathrm${GHz}), the system nearly stays in its ground state at such a low temperature. To see the coupling between the two modes of the resonator, we measured the transmission spectra by applying a weak probe tone with its frequency scanning across the lowest transition energies $\tilde{\omega}_{j}$ with $j=1,3$. The spectra are fitted with the numerically calculated excitation energies of the system Hamiltonian $H_{\mathrm{s}}$. The parameters are determined to be
${\omega}_{1}/2\pi=$ $\ 5.0/\sqrt{1-\beta_{1}\cos(\theta)|\delta{\mathrm{\Phi}}_{\mathrm{ext}}|}~\mathrm{GHz},$ (5a)
${\omega}_{2}/2\pi=$ $\ 9.7/\sqrt{1-\beta_{2}\cos(\theta)|\delta{\mathrm{\Phi}}_{\mathrm{ext}}|}~\mathrm{GHz},$ (5b)
$g_{1}/2\pi=$ $\ 2.815~\mathrm{GHz},$ (5c)
$g_{2}/2\pi=$ $\ 2.180~\mathrm{GHz},$ (5d)
with $\beta_{1}=0.775~\mathrm{(\Phi_{0})^{-1}}$ and $\beta_{2}=0.919~\mathrm{(\Phi_{0})^{-1}}$. The coupling strengths
${g_{1}}/{\omega_{1}}=$ $\ 0.563>0.1,$ (6a)
${g_{2}}/{\omega_{2}}=$ $\ 0.225>0.1$ (6b)
at the optimal point (${\delta\mathrm{\Phi}}_{\mathrm{ext}}=0$) provide evidence our system is in the multimode ultrastrong coupling regime. Numerical estimates of coupling strengths from circuit quantization qualitatively agree with the fitted values above (see Supplementary Information). Parameters from the fit are used as input for density matrix calculations which are then compared with the measured spectra. Theoretical spectra throughout this work are obtained by using a generalized master equation Petruccione2002Theory ; Settineri2018Dissipation , where the interaction with the environment takes into account the hybridization of the subsystems due to the ultrastrong coupling (see Methods for details).
Figure 2 reports the measured (left) and calculated (right) transmission spectra versus the flux offset. The calculated spectra are proportional to the time derivative of the expectation value of the two-mode field momentum operator:
${X}=i({a}_{1}-{a}_{1}^{\dagger})+i\sqrt{\omega_{2}/\omega_{1}}({a}_{2}-{a}_{2}^{\dagger})$ (7)
(see Supplemeantary Information for details). Figure 2a shows the lowest energy transition $\tilde{\omega}_{1}$ corresponding to the excitation of the state
$|E_{1}\rangle\equiv|\widetilde{1,0}\rangle\simeq|1,0,g\rangle.$ (8)
@@
Figure 2b shows the transitions at $\tilde{\omega}_{2}$ and $\tilde{\omega}_{3}$. When the two-photon resonance frequency of the $n=1$ mode of the resonator gets across the single-photon resonance frequency of the $n=2$ mode ($\omega_{2}\simeq 2\omega_{1}$), a clear avoided crossing, induced by the effective interaction in Eq. (3), is observed in Fig. 2b, indicating that our system enters the single-photon strong nonlinear coupling regime Bouwmeester2006PRL $g_{\mathrm{eff}}>\kappa_{\mathrm{tot}}$, where ${\kappa}_{\mathrm{tot}}$ is the total loss rate of the resonator. The nonlinear coupling strength $g_{\mathrm{eff}}/2\pi$ extracted from the fit is 59 MHz (half the separation between the double peaks at the minimum splitting), which qualitatively agrees with the perturbative result given by Eq. 4 (47 MHz). While for flux offsets outside the two-modes resonance condition, the system eigenstates are approximately $|2,0,g\rangle$ and $|0,1,g\rangle$; at the anticrossing they are in good approximation their symmetric and antisymmetric superpositions
$|\psi_{\pm}\rangle=|2,0,g\rangle\pm|0,1,g\rangle.$ (9)
Note that outside the avoided-level crossing in Fig. 2b, the spectral line corresponding to the excitation of the two-photon state $|2,0,g\rangle$ is poorly excited by the weak probe tone. Two-photon contributions get significantly excited only owing to their hybridization with one-photon states at the avoided crossing. The emission of photon pairs under very weak excitation is maximized in this region (see Supplementary Fig. 3). Theoretical calculations accurately reproduce all the spectra in Fig. 2b, thus confirming our interpretation of the data. A peculiar feature of this extreme quantum nonlinear regime is that linear spectra in Fig. 2b are able to evidence quantum nonlinear optical processes.